The integral of re-ar is easier to evaluate when n is odd (a) Evaluate Lºre-*z de (b) Evaluate the indefinite integral of replease write out all work
Answer #1

@ an odd ben)=- ne-a langt E-ne-ant = – fin) . xe-an du since ben) = re-ane is I (bende blonde function as - joo (sens - blas